OpenStudy (anonymous):

I need help with my last question!!!! I will give medal!!!! Part A: The area of a square is (16x^2 -8x+1) square units. Determine the length of each side of the square by factoring the area expression completely. Show your work. Part B: The area of a rectangle is (81x^2 -4y^2) square units. Determine the dimensions of the rectangle by factoring the area expression completely. Show your work. Part C: The volume of a rectangular box is (x^3 -9x^2 -4x+36) cubic units. Determine the dimensions of the rectangular box by factoring the volume expression completely. Show your work.

OpenStudy (anonymous):

\[x^2\left( x-9 \right)-4\left( x-9 \right)=0,\left( x-9 \right)\left( x^2-4 \right)=0\] \[\left( x-9 \right)\left( x+4 \right)\left( x-2 \right)=0\]

OpenStudy (anonymous):

correction write x+2 in place of x+4

OpenStudy (anonymous):

if x is positive then dimensions are x+2,x-2,x-9 don't write =0 just write = at every step.